OpenStudy (anonymous):

What method do you use to solve (8n+5) (n-4)=0?

OpenStudy (anonymous):

We use the method where multiplying anything by 0 nets you zero. So 8n+5=0, and n-4=0. Voila.

OpenStudy (anonymous):

If the product of the two expressions is 0, then one of them must equal zero. 8n+5 = 0 n - 4 = 0